OpenStudy (anonymous):

I am doing integrals by substitution, U=lnx, dx=(du/(1/x) please tell me if there is a better way to plot that, or if I messing up something.

OpenStudy (agreene):

what is your original integral?

OpenStudy (anonymous):

is \[\int\limits_{}^{}(lnx)^6\div x dx\]

OpenStudy (agreene):

Let u= ln x du = 1/x dx That gives us: \[\int u^6 du = \frac17 u^7 + C\] plug back in for u: \[=\frac{(\ln x)^7}{7}+C\]

OpenStudy (anonymous):

Thanks man

OpenStudy (agreene):

no problem--it looked like you were basically there, you just havent gotten the hang of substitution all the way yet :P